OpenStudy (anonymous):

3x^2 - 16 -12 factor

OpenStudy (anonymous):

\[(3x + 2 )(x - 6)\]

OpenStudy (anonymous):

thank you for the answer but how do you get it?

OpenStudy (anonymous):

Well you need to know how to factor.

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

multiply the coefficient of x by 12 3 x -12 = -36

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i got stuck there

OpenStudy (anonymous):

Or you can use this equation\[x = -b \pm \sqrt{b ^{2} - 4ac} \div 2a\]

OpenStudy (anonymous):

no thank you

OpenStudy (anonymous):

there is an easier way

OpenStudy (anonymous):

what are the factors of -36 that = -16

OpenStudy (anonymous):

\[ax ^{2} + bx + c\] is the original equation to plug in numbers.

OpenStudy (anonymous):

i don't know that why i am stuck

OpenStudy (anonymous):

its already plugged in...im confused

OpenStudy (anonymous):

So in that \[3x ^{2}- 16x -12\] a = 3 b = -16 c = -12

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i just need a common factor that adds to -16 but multiplies to -36

OpenStudy (anonymous):

Here: http://www.tc3.edu/instruct/sbrown/ti83/quadrat.htm. This is a good program to plug into your calculator. Quadratics are something that you will always deal with and this will make it easier.

OpenStudy (anonymous):

i don't need a calculater

OpenStudy (anonymous):

I'm just saying for the future when you get quadratic equations like \[47x ^{2} -183x + 95\]

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

or 3454546565656434x^2+54496054x+0.434343883948=0

OpenStudy (anonymous):

what

OpenStudy (anonymous):

or even 5894854985985x^2+8485x+0.3434=0

OpenStudy (anonymous):

if the method i was trying dosent work, and it looks like the factors are not working then use the formula Doc gave you because that will work every time.

OpenStudy (anonymous):

i kow how to factor!!!

OpenStudy (anonymous):